# Measurement of the power spectrum turnover scale from the cross-correlation between CMB lensing and Quaia

### v2

@@ Abstract
We use the projected clustering of quasars in the Gaia-unWISE quasar catalog, Quaia, and its cross-correlation with CMB lensing data from Planck, to measure the large-scale turnover of the matter power spectrum, associated with the size of the horizon at the epoch of matter-radiation equality. The turnover is detected with a significance of between $2.3$ and $3.1\sigma$, depending on the method used to quantify it. From this measurement, the equality scale is determined at the $\sim 20\%$ level. Using the turnover scale as a standard ruler alone (discarding information from the large-scale curvature of the power spectrum), in combination with supernova data through an inverse distance ladder approach, we measure the current expansion rate to be $H_{0}=62.7\pm 17.2\,{\rm km}\,{\rm s}^{-1}\,{\rm Mpc}^{-1}$. The addition of information coming from the power spectrum curvature approximately halves the standard ruler uncertainty. Our measurement in combination with calibrated supernovae from Pantheon$+$ and SH0ES constrains the CMB temperature to be $T_{\rm CMB}=3.10^{+0.48}_{-0.36}\,{\rm K}$, independently of CMB data. Alternatively, assuming the value of $T_{\rm CMB}$ from COBE-FIRAS, we can constrain the effective number of relativistic species in the early Universe to be $N_{\rm eff}=3.0^{+5.8}_{-2.9}$.
@@ 1 Introduction
The study of the Universe’s large-scale structure (LSS) aims to extract cosmological information from the distribution of matter inhomogeneities on different scales, and their evolution in time. Key to these studies is the linear matter power spectrum $P_{L}(k)$: the Fourier-space variance of the matter overdensity at early times [^1] It is common to work with $P_{L}(k)$ at $z=0$ by simply evolving it using linear theory.. The form of $P_{L}(k)$ is fully determined by a small number of characteristic scales, associated with specific physical phenomena affecting the growth of structure. On small scales, neutrino free streaming causes power suppression above a characteristic wavenumber $k_{\rm fs}$, related to the neutrino mass (Lesgourgues and Pastor, 2006; Bird et al., 2012). Another key scale is the sound horizon of the baryon-photon plasma at the time of decoupling $r_{d}$, which sets the frequency of the baryon acoustic oscillations (BAO, Eisenstein and Hu, 1998), a central observable for LSS that brought the field into the regime of precision cosmology (Eisenstein et al., 2005). Finally, the physical scale that gives rise to the most prominent feature in the matter power spectrum is the size of the horizon at the time of matter-radiation equality. Matter fluctuations on scales below this horizon were subject to a suppressed growth during the radiation-dominated era, whereas larger scales evolved largely unaffected into the matter epoch. This gives rise to a prominent feature in the matter power spectrum: a peak, or turnover, at a wavenumber close to the comoving horizon at the matter-radiation equality $k_{\rm eq}\sim 0.01\,{\rm Mpc}^{-1}$, below which $P_{L}(k)$ grows as $\propto k^{n_{s}}$ (with $n_{s}$ the scalar spectral index), and above which it decays as $\sim k^{n_{s}-4}$.
@@
The power spectrum turnover has significant value for cosmology. Much like the BAO scale, the turnover feature may be used as a standard ruler to place constraints on the distance-redshift relation. Furthermore, the dependence on cosmological parameters of the ruler (i.e. the size of the comoving horizon at matter-radiation equality) is highly complementary to that of other probes of the background expansion (e.g. supernovae and BAO), and thus can be used in combination with them to break parameter degeneracies and obtain complementary constraints on key parameters such as the late-time expansion rate $H_{0}$, the matter density parameter $\Omega_{m}$ and, potentially, the Dark Energy equation of state $w$. Interestingly, this may be achieved even without a clear detection of the turnover itself, relying instead on the broadband curvature in the power spectrum on smaller scales caused by the turnover. This is at the heart of several approaches recently employed in the literature, including full-shape power spectrum analyses (Philcox et al., 2021), the ShapeFit methodology (Brieden et al., 2021), and exploiting the curvature of the power spectrum of the Cosmic Microwave Background (CMB) lensing convergence (Philcox et al., 2022).
@@
However, in spite of its prominence, the power spectrum turnover has been surprisingly difficult to detect directly in galaxy survey data. The main reason for this is the comparably large scale of this feature. Subtended on the sky, the equality scale would correspond to separations of about 10 at $z\sim 0.5$ and 3 at $z=2$. Obtaining reliable measurements on such large scales from galaxy survey data is complicated for two reasons: first, samples covering very large volumes are required in order to reduce the statistical uncertainties sufficiently. Secondly, on these scales the observed galaxy overdensity can be dominated by observational systematics: the fluctuations in the observed number of galaxies caused by non-cosmological effects. These include Galactic contaminats, most prominently dust extinction and star contamination, as well as instrumental effects, such as variations in observing conditions leading to an inhomogeneous survey depth. For this reason, to our knowledge, only two attempts have been made at detecting the power spectrum turnover from galaxy survey data. An early tentative detection was made with the WiggleZ survey (Poole et al., 2013) and, more recently, the turnover was measured using quasars from the extended Baryon Oscillation Spectroscopic Survey, eBOSS (Bahr-Kalus et al., 2023). In both cases, these results were obtained using the 3D auto-correlation of galaxies, and thus significant effort had to be devoted to ensuring the robustness of the measurements against instrumental and survey systematics. Being able to make the same measurement using other LSS probes, with different sensitivity to these systematics, is therefore of vital importance.
@@ 2.1 Quaia
We divide the full sample into the same 2 redshift bins used in Alonso et al. (2023), corresponding to sources with redshifts above and below the median redshift of the sample $z_{Q}^{\rm med}=1.47$ (we will refer to these as the “High-$z$” and “Low-$z$” bins, respectively). The ability to select different redshift ranges is extremely useful in this case, as it allows us to quantify whether the projected turnover scale scales with redshift as expected. We will also study the results obtained from the full sample as an additional internal consistency test. A selection function $w(\hat{\bf n})$ was generated for each redshift bin characterising the fluctuations in the mean density of sources caused by spatially-varying systematics. This selection function is generated via Gaussian process reconstruction as described in Storey-Fisher et al. (2024). As described in Section 3.1.1, we deproject a number of systematic templates from the quasar ovedensity maps prior to power spectrum estimation. These include the templates used to create the selection functions.
@@ 2.2 Planck PR4 lensing map
We use CMB lensing maps reconstructed from the fourth Planck public data release (PR4), generated through the NPIPE pipeline (Planck Collaboration et al., 2020a), and presented in Carron et al. (2022a). In particular, we made use of the “Generalised Minimum Variance” (GMV) convergence map, which implements a joint inverse-variance Wiener filtering of the temperature and polarisation data accounting for inhomogeneous noise, achieving a $\sim 20\%$ better signal-to-noise ratio than previous releases. Although the cosmological analysis of the CMB lensing auto-spectrum used the multipole range $\ell\geq 8$ (Carron et al., 2022a), we compute the cross-correlation with Quaia down to $\ell=2$ (and show that our results are largely insensitive to this choice). The choice $\ell\geq 8$ for the lensing power spectrum analysis stems from concerns about the reliability of the subtracted mean field Planck Collaboration et al. (2020b), which is less relevant for cross-correlations both with external data (Krolewski et al., 2024) and other Planck CMB observables (Carron et al., 2022b).
@@ 3.1.2 Transfer functions
In a similar manner, we must account for the effects of mode loss in the lensing reconstruction maps. The presence of a sky mask leads to a mis-normalisation of the reconstructed lensing map. Since this effect is inhomogeneous, its impact on the final power spectrum depends on the sky mask of the fields being cross-correlated (Benoit-Lévy et al., 2013; Planck Collaboration et al., 2020b; Carron, 2023; Krolewski et al., 2024). We estimate the transfer function associated with this effect making use of the lensing reconstruction simulations made available with the PR4 lensing maps [^5] See https://github.com/carronj/planck_PR4_lensing., as described in Farren et al. (2024). For each realisation, three maps are generated: $\kappa_{\kappa}$, the input convergence field masked with the CMB lensing analsis mask, $\kappa_{g}$, the input convergence field masked with the Quaia selection function, and $\kappa_{\rm rec}$, the reconstructed convergence field (masked with the lensing mask). The transfer function is then:
@@ 3.2.1 Angular power spectra
where $p(z)$ is the redshift distribution of the galaxy sample under study, $H(z)$ is the expansion rate at redshift $z$, $H_{0}$ is its value today, $\Omega_{m}$ is the fractional energy density in non-relativistic species today, and $\chi_{*}$ is the distance to the last-scattering surface.
@@ 3.2.2 Modelling the power spectrum turnover
$P_{gg}(k,z)=b_{g}^{2}P_{mm}(k,z),\hskip 12.0ptP_{gm}(k,z)=b_{g}P_{mm}(k,z),$ (13)
where $b_{g}$ is the linear Galaxy bias.
@@ 3.2.3 Likelihood
where ${\bf d}$ is the data vector, $\vec{\theta}$ is the set of free parameters in our model, ${\bf t}$ is the theory prediction, and ${\sf C}$ is the covariance matrix [^8] We correct ${\sf C}^{-1}$ by the so-called “Hartlap factor” (Hartlap et al., 2007) to account for the bias in covariance matrices estimated from a finite number of simulations..
@@ 3.2.4 The projected turnover scale as a standard ruler
To interpret our measurement of the power spectrum turnover in a cosmological setting, we must connect the stretch parameter $\alpha$ with the cosmological parameters governing the background expansion of the Universe. As discussed in Section 3.2.1, a turnover in the matter power spectrum at a comoving scale $k_{\rm TO}$ would map onto a turnover in the angular power spectra at an angular scale $\ell_{\rm TO}\approx k_{\rm TO}d_{A}(\bar{z})$, where $d_{A}(\bar{z})$ is the comoving angular diameter distance at the median redshift of the sample under study. In what follows, we will only interpret our measurements in terms of a flat $\Lambda$CDM model, and therefore $d_{A}$ is the same as the radial comoving distance $\chi$. The model described in Sections 3.2.1 and 3.2.2 would predict this turnover at an angular scale $l$ given by $k_{\rm TO}^{\rm fid}\,d_{A}^{\rm fid}(\bar{z})/\alpha$, where the “fid” superscript denotes quantities estimated in the fiducial cosmological model assumed (the best-fit Planck $\Lambda$CDM parameters, in our case). We can therefore interpret our constraints on $\alpha$ as a measurement of the combination
@@ 4.1 Detection of the power spectrum turnover
Although the theoretical predictions are useful in guiding the eye, it is not immediately evident that the data shows a clear preference for a turnover in the power spectrum, compared with a plateau or a monotonically-decreasing trend. To quantitatively determine our confidence that the turnover is indeed detected, we can use the posterior distribution of $\beta$: since negative values of $\beta$ correspond to spectra with no turnover (with $\beta=0$ a plateau), the detection significance of the turnover can be determined from the probability $p(\beta<0)$. This marginalised posterior is shown in black in Fig 8. We find that, with our fiducial analysis choices, 99.83% of the MCMC samples correspond to positive $\beta$ values, corrsponding to a $3.1\sigma$ detection of a turnover in the power spectrum. The best-fit value of $\beta$ in our fiducial setup is $\beta=3.22$, steeper than the $\Lambda$CDM value $\beta=1$ (which is nevertheless compatible with our constraints). This result is robust against changes in this choices: repeating our analysis combining all Quaia sources into a single redshift bin (red dashed line), we find $p(\beta<0)=0.0035$ ($2.9\sigma$). Since the quasar auto-correlation is more sensitive to systematic contamination on large scales, it is also interesting to explore the constraints obtained from an analysis involving only the cross-correlations with CMB lensing. This results in a mild reduction of the detection significance $p(\beta<0)=0.0064$ ($2.7\sigma$ – dash-dotted line in Fig. 8), but still a clear preference for a turnover. Finally, as a large-scale effect, the detectability of the turnover may depend critically on the lowest multipoles used in the analysis. Since these may be more sensitive to inaccuracies in the covariance matrix or even the shape of the likelihood itself, we repeat the analysis using a more conservative large-scale cut, with $\ell_{\rm min}=10$ for the cross-correlations. The resulting posterior (dotted yellow line) still exhibits a clear preference for the presence of a turnover, with a similar confidence level ($p(\beta<0)=0.009$, or $2.6\sigma$).
@@
For a Bayesian analyst this is the end of the story – the result depends on the prior and for sensible priors on $\beta$ we get a significant, but not decisive detection. To satisfy the working class, we also carry out a frequentist test, which does not depend on any $\beta$ priors, as an alternative way to quantify this detection significance. We generate 1000 random realisations of the measured power spectra, drawn from a multivariate normal with a mean given by the best-fit theoretical prediction derived from our fiducial analysis, and the covariance matrix of our data. We then find the best-fit values of all free parameters in the model ($\alpha$, $\beta$, $\gamma$, and the power spectrum amplitudes) for each realisation, and count the number of realisations for which the best-fit $\beta$ is negative. Through this exercise, we find $p(\beta<0)=0.02$ (i.e. 20 realisations were found with $\beta<0$), corresponding to a $2.3\sigma$ detection. A similar exercise drawing realisations with a mean given by a theoretical prediction with $\beta=0$ (i.e. a limiting model with no turnover) found only $32$ out of 1000 realisations (i.e. $p=0.032$) with a best-fit value of $\beta$ above that found in our data.

### v3

@@ Abstract
We use the projected clustering of quasars in the Gaia-unWISE quasar catalog, Quaia, and its cross-correlation with CMB lensing data from Planck, to measure the large-scale turnover of the matter power spectrum, associated with the size of the horizon at the epoch of matter-radiation equality. The turnover is detected with a significance of between $2.3$ and $3.1\sigma$, depending on the method used to quantify it. From this measurement, the equality scale is determined at the $\sim 20\%$ level. Using the turnover scale as a standard ruler alone (suppressing information from the large-scale curvature of the power spectrum), in combination with supernova data through an inverse distance ladder approach, we measure the current expansion rate to be $H_{0}=62.7\pm 17.2\,{\rm km}\,{\rm s}^{-1}\,{\rm Mpc}^{-1}$. The addition of information coming from the power spectrum curvature approximately halves the standard ruler uncertainty. Our measurement in combination with calibrated supernovae from Pantheon$+$ and SH0ES constrains the CMB temperature to be $T_{\rm CMB}=3.10^{+0.48}_{-0.36}\,{\rm K}$, independently of CMB data. Alternatively, assuming the value of $T_{\rm CMB}$ from COBE-FIRAS, we can constrain the effective number of relativistic species in the early Universe to be $N_{\rm eff}=3.0^{+5.8}_{-2.9}$.
@@ 1 Introduction
The study of the Universe’s large-scale structure (LSS) aims to extract cosmological information from the distribution of matter inhomogeneities on different scales, and their evolution in time. Key to these studies is the linear matter power spectrum $P_{L}(k)$: the Fourier-space variance of the matter overdensity at early times [^1] It is common to work with $P_{L}(k)$ at $z=0$ by simply evolving it using linear theory.. The form of $P_{L}(k)$ is fully determined by a small number of characteristic scales, associated with specific physical phenomena affecting the growth of structure. On small scales, neutrino free streaming causes power suppression above a characteristic wavenumber $k_{\rm fs}$, related to the neutrino mass (Lesgourgues and Pastor, 2006; Bird et al., 2012). Another key scale is the sound horizon of the baryon-photon plasma at the time of decoupling $r_{d}$, which sets the frequency of the baryon acoustic oscillations (BAO, Eisenstein and Hu, 1998), a central observable for LSS that brought the field into the regime of precision cosmology (Eisenstein et al., 2005). Finally, the physical scale that gives rise to the most prominent feature in the matter power spectrum is the size of the horizon at the time of matter-radiation equality. Matter fluctuations on scales below this horizon were subject to a suppressed growth during the radiation-dominated era, whereas larger scales evolved largely unaffected into the matter epoch. This gives rise to a prominent feature in the matter power spectrum: a peak, or turnover, at a wavenumber close to the comoving horizon at the matter-radiation equality $k_{\rm eq}\sim 0.01\,{\rm Mpc}^{-1}$, below which $P_{L}(k)$ grows as $\propto k^{n_{s}}$ (with $n_{s}$ the scalar spectral index), and above which it decays as $\sim k^{n_{s}-4}$. Specifically, this change of slope, from positive on large scales, to negative on small scales, and the scale at which it takes place, is what we will refer to as “the turnover” in what follows.
@@
The power spectrum turnover has significant value for cosmology. Much like the BAO scale, the turnover feature may be used as a standard ruler to place constraints on the distance-redshift relation. Furthermore, the dependence on cosmological parameters of the ruler (i.e. the size of the comoving horizon at matter-radiation equality) is highly complementary to that of other probes of the background expansion (e.g. supernovae and BAO), and thus can be used in combination with them to break parameter degeneracies and obtain complementary constraints on key parameters such as the late-time expansion rate $H_{0}$, the matter density parameter $\Omega_{m}$ and, potentially, the Dark Energy equation of state $w$. Interestingly, this may be achieved even without a clear detection of the turnover itself, relying instead on the broadband curvature in the power spectrum on smaller scales caused by the turnover. Several approaches recently employed in the literature, including full-shape power spectrum analyses (Philcox et al., 2021), the ShapeFit methodology (Brieden et al., 2021), and exploiting the curvature of the power spectrum of the Cosmic Microwave Background (CMB) lensing convergence (Philcox et al., 2022), benefit from this smaller-scale curvature.
@@
However, in spite of its prominence, the power spectrum turnover has been surprisingly difficult to detect directly in galaxy survey data. The main reason for this is the comparably large scale of this change in slope. Subtended on the sky, the equality scale would correspond to separations of about 10 at $z\sim 0.5$ and 3 at $z=2$. Obtaining reliable measurements on such large scales from galaxy survey data is complicated for two reasons: first, samples covering very large volumes are required in order to reduce the statistical uncertainties sufficiently. Secondly, on these scales the observed galaxy overdensity can be dominated by observational systematics: the fluctuations in the observed number of galaxies caused by non-cosmological effects. These include Galactic contaminants, most prominently dust extinction and star contamination, as well as instrumental effects, such as variations in observing conditions leading to an inhomogeneous survey depth. In turn, measuring the small-scale slope of the power spectrum is significantly simpler. For this reason, to our knowledge, only two attempts have been made at detecting the power spectrum turnover from galaxy survey data. An early tentative detection was made with the WiggleZ survey (Poole et al., 2013) and, more recently, the turnover was measured using quasars from the extended Baryon Oscillation Spectroscopic Survey, eBOSS (Bahr-Kalus et al., 2023). In both cases, these results were obtained using the 3D auto-correlation of galaxies, and thus significant effort had to be devoted to ensuring the robustness of the measurements against instrumental and survey systematics. Being able to make the same measurement using other LSS probes, with different sensitivity to these systematics, is therefore of vital importance.
@@ 2.1 Quaia
We divide the full sample, spanning the redshift range $0<z\lesssim 4$, into the same 2 redshift bins used in Alonso et al. (2023), corresponding to sources with redshifts above and below the median redshift of the sample $z_{Q}^{\rm med}=1.47$ (we will refer to these as the “High-$z$” and “Low-$z$” bins, respectively). The ability to select different redshift ranges is extremely useful in this case, as it allows us to quantify whether the projected turnover scale scales with redshift as expected. We will also study the results obtained from the full sample as an additional internal consistency test. A selection function $w(\hat{\bf n})$ was generated for each redshift bin characterising the fluctuations in the mean density of sources caused by spatially-varying systematics. This selection function is generated via Gaussian process reconstruction as described in Storey-Fisher et al. (2024). As described in Section 3.1.1, we deproject a number of systematic templates from the quasar ovedensity maps prior to power spectrum estimation. These include the templates used to create the selection functions.
@@ 2.2 Planck PR4 lensing map
We use CMB lensing maps reconstructed from the fourth Planck public data release (PR4), generated through the NPIPE pipeline (Planck Collaboration et al., 2020a), and presented in Carron et al. (2022a). In particular, we made use of the “Generalised Minimum Variance” (GMV) convergence map, which implements a joint inverse-variance Wiener filtering of the temperature and polarisation data accounting for inhomogeneous noise, achieving a $\sim 20\%$ better signal-to-noise ratio than previous releases. Although the cosmological analysis of the CMB lensing auto-spectrum used the multipole range $\ell\geq 8$ (Carron et al., 2022a), we compute the cross-correlation with Quaia down to $\ell=2$. Our fiducial analysis will use scales $\ell\geq 4$, and we will show that the results are insensitive to this choice. The choice $\ell\geq 8$ for the Planck lensing power spectrum analysis stems from concerns about the reliability of the subtracted mean field Planck Collaboration et al. (2020b), which is less relevant for cross-correlations both with external data (Krolewski et al., 2024) and other Planck CMB observables (Carron et al., 2022b).
@@ 3.1.2 Transfer functions
In a similar manner, we must account for the effects of mode loss in the lensing reconstruction maps. The presence of a sky mask leads to a mis-normalisation of the reconstructed lensing map. Since this effect is inhomogeneous, its impact on the final power spectrum depends on the sky mask of the fields being cross-correlated (Benoit-Lévy et al., 2013; Planck Collaboration et al., 2020b; Carron, 2023; Krolewski et al., 2024). We estimate the transfer function associated with this effect making use of the lensing reconstruction simulations made available with the PR4 lensing maps [^5] See https://github.com/carronj/planck_PR4_lensing., as described in Farren et al. (2024). For each realisation, three maps are generated: $\kappa_{\kappa}$, the input convergence field masked with the CMB lensing analysis mask, $\kappa_{g}$, the input convergence field masked with the Quaia selection function, and $\kappa_{\rm rec}$, the reconstructed convergence field (masked with the lensing mask). The transfer function is then:
@@ 3.2.1 Angular power spectra
where $p(z)$ is the redshift distribution of the galaxy sample under study, $H(z)$ is the expansion rate at redshift $z$, $H_{0}$ is its value today, $\Omega_{m}$ is the fractional energy density in non-relativistic species today, and $\chi_{*}$ is the distance to the last-scattering surface. Note that we will assume a flat $\Lambda$CDM cosmology in our analysis, in which case the radial comoving distance $\chi(z)$ is equivalent to the comoving angular diameter distance $d_{A}(z)$, which we will refer to in Section 3.2.4.
@@ 3.2.2 Modelling the power spectrum turnover
$P_{gg}(k,z)=b_{g}(z)^{2}P_{mm}(k,z),\hskip 6.0ptP_{gm}(k,z)=b_{g}(z)P_{mm}(k,z),$ (13)
where $b_{g}(z)$ is the redshift-dependent linear Galaxy bias.
@@ 3.2.3 Likelihood
where ${\bf d}$ is the data vector, $\vec{\theta}$ is the set of free parameters in our model, ${\bf t}$ is the theory prediction, ${\sf C}$ is the covariance matrix [^8] We correct ${\sf C}^{-1}$ by the so-called “Hartlap factor” (Hartlap et al., 2007) to account for the bias in covariance matrices estimated from a finite number of simulations., and $K$ is an arbitrary normalisation constant.
@@ 3.2.4 The projected turnover scale as a standard ruler
To interpret our measurement of the power spectrum turnover in a cosmological setting, we must connect the stretch parameter $\alpha$ with the cosmological parameters governing the background expansion of the Universe. As discussed in Section 3.2.1, a turnover in the matter power spectrum at a comoving scale $k_{\rm TO}$ would map onto a turnover in the angular power spectra at an angular scale $\ell_{\rm TO}\approx k_{\rm TO}d_{A}(\bar{z})$, where $d_{A}(\bar{z})$ is the comoving angular diameter distance at the median redshift of the sample under study. The model described in Sections 3.2.1 and 3.2.2 would predict this turnover at an angular scale $l$ given by $k_{\rm TO}^{\rm fid}\,d_{A}^{\rm fid}(\bar{z})/\alpha$, where the “fid” superscript denotes quantities estimated in the fiducial cosmological model assumed (the best-fit Planck $\Lambda$CDM parameters, in our case). We can therefore interpret our constraints on $\alpha$ as a measurement of the combination
@@ 4.1 Detection of the power spectrum turnover
Although the theoretical predictions are useful in guiding the eye, it is not immediately evident that the data shows a clear preference for a turnover in the power spectrum, compared with a plateau or a monotonically-decreasing trend. To quantitatively determine our confidence that the turnover is indeed detected, we can use the posterior distribution of $\beta$: since negative values of $\beta$ correspond to spectra with no turnover (with $\beta=0$ a plateau), the detection significance of the turnover can be determined from the probability $p(\beta<0)$. This marginalised posterior is shown in black in Fig 8. We find that, with our fiducial analysis choices, 99.83% of the MCMC samples correspond to positive $\beta$ values, corresponding to a $3.1\sigma$ detection of a turnover in the power spectrum. The best-fit value of $\beta$ in our fiducial setup is $\beta=3.22$, steeper than the $\Lambda$CDM value $\beta=1$ (which is nevertheless compatible with our constraints). This result is robust against changes in these choices: repeating our analysis combining all Quaia sources into a single redshift bin (red dashed line), we find $p(\beta<0)=0.0035$ ($2.9\sigma$). Since the quasar auto-correlation is more sensitive to systematic contamination on large scales, it is also interesting to explore the constraints obtained from an analysis involving only the cross-correlations with CMB lensing. This results in a mild reduction of the detection significance $p(\beta<0)=0.0064$ ($2.7\sigma$ – dash-dotted line in Fig. 8), but still a clear preference for a turnover. Finally, as a large-scale effect, the detectability of the turnover may depend critically on the lowest multipoles used in the analysis. Since these may be more sensitive to inaccuracies in the covariance matrix or even the shape of the likelihood itself, we repeat the analysis using a more conservative large-scale cut, with $\ell_{\rm min}=10$ for the cross-correlations. The resulting posterior (dotted yellow line) still exhibits a clear preference for the presence of a turnover, with a similar confidence level ($p(\beta<0)=0.009$, or $2.6\sigma$).
@@
For a Bayesian analyst this is the end of the story – the result depends on the prior and for sensible priors on $\beta$ we get a significant, but not decisive detection. To satisfy the working class, we also carry out a frequentist test, which does not depend on any $\beta$ priors, as an alternative way to quantify this detection significance. We generate 1000 random realisations of the measured power spectra, drawn from a multivariate normal with a mean given by the best-fit theoretical prediction derived from our fiducial analysis, and the covariance matrix of our data. We then find the best-fit values of all free parameters in the model ($\alpha$, $\beta$, $\gamma$, and the power spectrum amplitudes) for each realisation, and count the number of realisations for which the best-fit $\beta$ is negative. Through this exercise, we find $p(\beta<0)=0.02$ (i.e. 20 realisations were found with $\beta<0$), corresponding to a $2.3\sigma$ detection. A similar exercise drawing realisations with a mean given by a theoretical prediction with $\beta=0$ (i.e. a limiting model with no turnover) found only $32$ out of 1000 realisations (i.e. $p=0.032$) with a best-fit value of $\beta$ above that found in our data. In other words, the limiting case of a power spectrum with no clear turnover, is disfavoured at the $2.15\sigma$ level.